The value of a used car is now 5,600 euros. What was the value of the car three years ago when it has dropped 16% annually?


1
Expert's answer
2021-06-08T14:45:36-0400

Let "A_0="  the value of the car three years ago.

Then


"A_3=A_0(1-0.16)^3=5600"

"A_0=\dfrac{5600}{(0.84)^3}=9448.22"

The value of the car three years ago was 9448.22 euros.